The acceleration of an object varies directly as the force acting on it. If a force of 240 newtons causes an acceleration of 150 m/sē, what force will cause an acceleration of 100 m/s2?
Answer by stanbon(57337)

The acceleration of an object varies directly as the force acting on it. If a force of 240 newtons causes an acceleration of 150 m/sē, what force will cause an acceleration of 100 m/s2?
---
a = kf
Solve for "k" using "a force of 240 newtons causes an acceleration of 150 m/sē"
---
150 = k*240
k = 0.625
----
What force will cause an acceleration of 100 m/s2?
Equation:
a = 0.625f
100 = 0.625*f
f = 100/0.625
f = 160 newtons
======================
cheers,
Stan H.